Use the following linear regression equation to answer the questions.

x1 = 1.4 + 3.1x2 – 8.2x3 + 2.1x4

Suppose x3 and x4 were held at fixed but arbitrary values and x2 increased by 1 unit. What would be the corresponding change in x1?


Suppose x2 increased by 2 units. What would be the expected change in x1?


Suppose x2 decreased by 4 units. What would be the expected change in x1?


(e) Suppose that n = 13 data points were used to construct the given regression equation and that the standard error for the coefficient of x2 is 0.461. Construct a 95% confidence interval for the coefficient of x2. (Use 2 decimal places.)

lower limit
upper limit


(f) Using the information of part (e) and level of significance 1%, test the claim that the coefficient of x2 is different from zero. (Use 2 decimal places.)

t
t critical ±
Answer #1

Answer: The given Equation is Xi = 1.44 3.1724273+2-124 Obtain the Constant term and obtain the table. Constant 1.47 3.1 X-Co(2 LE I-CI = 1-0.95 -0.05 df = n-4-13-4-9 Gitical value of t to.o5 씰 to.025 A *2.571 - 95% Confidence interval is, C.I B2 + t